### v3

# On cyclotomic matrices involving Gauss sums over finite fields

Abstract.
Inspired by the works of L. Carlitz and Z.-W. Sun on cyclotomic matrices, in this paper, we investigate certain cyclotomic matrices involving Gauss sums over finite fields, which can be viewed as finite field analogues of certain matrices related to the Gamma function.
For example, let $q=p^{n}$ be an odd prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\zeta_{p}=e^{2\pi{\bf i}/p}$ and let $\chi$ be a generator of the group of all mutiplicative characters of the finite field $\mathbb{F}_{q}$. For the Gauss sum
$G_{q}(\chi^{r})=\sum_{x\in\mathbb{F}_{q}}\chi^{r}(x)\zeta_{p}^{{\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}(x)},$
we prove that
$\det\left[G_{q}(\chi^{2i+2j})\right]_{0\leq i,j\leq(q-3)/2}=(-1)^{\alpha_{p}}\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}2^{\frac{p^{n-1}-1}{2}},$
where
$\alpha_{p}=\begin{cases}1&\mbox{if}\ n\equiv 1\ ({\rm{mod}}\ 2),\\ (p^{2}+7)/8&\mbox{if}\ n\equiv 0\ ({\rm{mod}}\ 2).\end{cases}$
Key words and phrases:
Gauss Sums, Finite Fields, Cyclotomic Matrices, Determinants.
2020 Mathematics Subject Classification. Primary 11L05, 15A15; Secondary 11R18, 12E20.
This work was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12201291).
1. Introduction
1.1. Notations
Let $q=p^{n}$ be a prime power with $p$ prime and $n\in\mathbb{Z}^{+}$ and let $\mathbb{F}_{q}$ be the finite field of $q$ elements. Let $\mathbb{F}_{q}^{\times}$ be the cyclic group of all nonzero elements of $\mathbb{F}_{q}$, and let $\widehat{\mathbb{F}_{q}^{\times}}$ be the cyclic group of all multiplicative characters of $\mathbb{F}_{q}$.
Throughout this paper, for any $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, we define $\psi(0)=0$. Also, we use the symbol $\varepsilon$ to denote the trivial multiplicative character of $\mathbb{F}_{q}$, i.e.,
$\varepsilon(x)=\begin{cases}1&\mbox{if}\ x\in\mathbb{F}_{q}^{\times},\\ 0&\mbox{if}\ x=0.\end{cases}$
Let $\zeta_{p}=e^{2\pi{\bf i}/p}$ and let ${\rm Tr}:={\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}$ be the trace map from $\mathbb{F}_{q}$ to $\mathbb{F}_{p}$. For any $x\in\mathbb{F}_{q}$, we have
${\rm Tr}(x)=\sum_{j=0}^{n-1}x^{p^{j}}\in\mathbb{F}_{p}.$
Let $\zeta_{q-1}=e^{2\pi{\bf i}/(q-1)}$. For any $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, the Gauss sum $G_{q}(\psi)$ over $\mathbb{F}_{q}$ is defined by
(1.1) $G_{q}(\psi)=\sum_{x\in\mathbb{F}_{q}}\psi(x)\zeta_{p}^{{\rm Tr}(x)}\in\mathbb{Q}(\zeta_{q-1},\zeta_{p}).$
1.2. Backgroud and Motivations
Gauss sums have been extensively studied and have many significant applications in both number theory and combinatorics. For example, In 1805 Gauss first determined the explicit values of quadratic Gauss sums over $\mathbb{F}_{p}$, which states that
(1.2) $\sum_{x\in\mathbb{F}_{p}}\left(\frac{x}{p}\right)\zeta_{p}^{x}=\begin{cases}p^{1/2}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ {\bf i}p^{1/2}&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4),\end{cases}$
where $(\frac{\cdot}{p})$ is the Legendre symbol, i.e., the unique quadratic multiplicative character of $\mathbb{F}_{p}$.
Also, the Hasse-Davenport lifting formula says that for any positive integer $n$ we have
(1.3) $G_{q}(\psi_{(n)})=(-1)^{n-1}G_{p}(\psi)^{n},$
where $q=p^{n}$ and $\psi_{(n)}=\psi\circ{\rm Norm}_{\mathbb{F}_{q}/\mathbb{F}_{p}}$ is a multiplicative character of $\mathbb{F}_{q}$.
Using the Hasse-Davenport lifting formula (1.3), one can generalize (1.2) to $\mathbb{F}_{q}$. In fact, when $p$ is an odd prime, let $\phi$ be the unique quadratic multiplicative character of $\mathbb{F}_{q}$, i.e.,
$\phi(x)=\begin{cases}0&\mbox{if}\ x=0,\\ 1&\mbox{if}\ x\ \text{is a nonzero square},\\ -1&\mbox{otherwise}.\end{cases}$
Then it is known that
$G_{q}(\phi)=\sum_{x\in\mathbb{F}_{q}}\phi(x)\zeta_{p}^{{\rm Tr}(x)}=\begin{cases}(-1)^{n-1}q^{1/2}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ (-1)^{n-1}{\bf i}^{n}q^{1/2}&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4).\end{cases}$
Now we turn to Carlitz’s cyclotomic matrices. Carlitz first studied the arithmetic properties of cyclotomic matrices involving multiplicative characters of $\mathbb{F}_{p}$. Let $\psi\in\widehat{\mathbb{F}_{p}^{\times}}$ be a nontrivial character with ${\rm ord}(\psi)=f$, where
${\rm ord}(\psi)=\min\{r\in\mathbb{Z}^{+}:\ \psi^{r}=\varepsilon\}.$
Let $p$ be an odd prime. Carlitz [3, Theorem 5] considered the cyclotomic matrix
(1.4) $C_{p}(\psi):=\left[\psi(i+j)\right]_{1\leq i,j\leq p-1},$
and determined the explicit value of $\det C_{p}(\psi)$, which states that
(1.5) $\det C_{p}(\psi)=\begin{cases}(-1)^{(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\nmid f,\\ (-1)^{(p-1)/f}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=1,\\ (-1)^{(f+2)(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=-1.\end{cases}$
Along this line, Chapman further studied some variants of $C_{p}(\psi)$. For instance, Chapman [4] consider the matrix
$V_{p}=\left[\left(\frac{i+j-1}{p}\right)\right]_{1\leq i,j\leq(p-1)/2}.$
Observing that
$\left(\frac{p+1}{2}-i\right)+\left(\frac{p+1}{2}-j\right)\equiv-(i+j-1)\ ({\rm{mod}}\ p),$
one can verify that
$\det V_{p}=\det\left[\left(\frac{i+j-1}{p}\right)\right]_{1\leq i,j\leq(p-1)/2}=\left(\frac{-1}{p}\right)\det\left[\left(\frac{i+j}{p}\right)\right]_{1\leq i,j\leq(p-1)/2}.$
Although the structures of $C_{p}(\phi)$ and $V_{p}$ are very similar, the calculation of $\det V_{p}$ is much more complicated than that of $\det C_{p}(\phi)$. Surprisingly, $\det V_{p}$ is closely related to the real quadratic field $\mathbb{Q}(\sqrt{p})$. In fact, let $\varepsilon_{p}>1$ and $h_{p}$ be the fundamental unit and the class number of $\mathbb{Q}(\sqrt{p})$ respectively and write
$\varepsilon_{p}^{h_{p}}=a_{p}+b_{p}\sqrt{p}\ (a_{b},b_{p}\in\mathbb{Q}).$
Chapman [4] proved that
(1.6) $\det V_{p}=\begin{cases}(-1)^{(p-1)/4}2^{(p-1)/2}b_{p}&\mbox{if}\ p\equiv 1\ ({\rm{mod}}\ 4),\\ 0&\mbox{if}\ p\equiv 3\ ({\rm{mod}}\ 4).\end{cases}$
After introducing the above relevant research results, we now describe our research motivations.
For any complex number $x$ with Re$(x)>0$, the Gamma function is defined by
[…]
For the Gamma function, the well known Gauss multiplication formula (see [1, Theorem 1.5.2]) says that for any $m\in\mathbb{Z}^{+}$ and any complex number $x$ we have
(1.8) $(2\pi)^{(m-1)/2}\Gamma(mx)=m^{(mx-1/2)}\prod_{j=0}^{m-1}\Gamma\left(x+\frac{j}{m}\right).$
The finite field counterpart of (1.8) is the Hasse-Davenport product formula (see [2, Theorem 11.3.5]), which states that for any $\rho\in\widehat{\mathbb{F}_{q}^{\times}}$ with ${\rm ord}(\rho)=m$ and any $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, we have
(1.9) $\prod_{0\leq a\leq m-1}G_{q}(\psi\rho^{a})=-\psi^{-m}(m)G_{q}(\psi^{m})\prod_{0\leq a\leq m-1}G_{q}(\rho^{a}).$
In view of (1.1), (1.7), (1.8) and (1.9), the function
$G_{q}(\cdot):\ \widehat{\mathbb{F}_{q}^{\times}}\rightarrow\mathbb{Q}(\zeta_{q-1},\zeta_{p})$
by sending the character $\psi$ to the Gauss sum $G_{q}(\psi)$, is indeed a finite field analogue of the Gamma function $\Gamma(\cdot)$. Readers may refer to [6] for the detailed introduction on this topic.
Throughout the remaining part of this paper, we let $\chi$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Now motivated by the structure of Carlitz’s cyclotomic matrix $C_{p}(\psi)$ defined by (1.4), it is natural to consider the matrix
(1.10) $A_{q}(1):=\left[G_{q}(\chi^{i+j})\right]_{0\leq i,j\leq q-2}.$
On the other hand, Sun [11] studied the matrix
$S_{p}=\left[\left(\frac{i^{2}+j^{2}}{p}\right)\right]_{1\leq i,j\leq(p-1)/2},$
which involves the nonzero squares of $\mathbb{F}_{p}$.
If we treat $S_{p}$ as a matrix over $\mathbb{F}_{p}$, then Sun [11, Theorem 1.2] proved that $-\det S_{p}$ is always a nonzero square of $\mathbb{F}_{p}$.
[…]
In the isomorphism $\mathbb{F}_{q}^{\times}\cong\widehat{\mathbb{F}_{q}^{\times}}$, the nonzero squares of $\mathbb{F}_{q}$ correspond to the even characters $\chi^{2i}$ $(i=0,1,\cdots,(q-3)/2)$. Thus, inspired by Sun’s matrix $S_{p}$ and the construction of $A_{q}(1)$, it is also natural to consider the matrix
(1.11) $A_{q}(2):=\left[G_{q}(\chi^{2i+2j})\right]_{0\leq i,j\leq(q-3)/2}.$
Motivated by the above results, in general, for any positive integer $k\mid q-1$, we define
(1.12) $A_{q}(k):=\left[G_{q}(\chi^{ki+kj})\right]_{0\leq i,j\leq(q-1-k)/k},$
and
(1.13) $B_{q}(k):=\left[G_{q}(\chi^{ki+kj})^{-1}\right]_{0\leq i,j\leq(q-1-k)/k},$
which concern the $k$th multiplicative characters of $\mathbb{F}_{q}$.
1.3. Main Results
We now state our main results of this paper.
Theorem 1.1.
Let $q=p^{n}$ be a prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\chi$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Let $k\mid q-1$ be a positive integer and let $m=(q-1)/k$. Then the following hold.
(i) $\det A_{q}(k)\in\mathbb{Z}$, $\det B_{q}(k)\in\mathbb{Q}$ and both are independent of the choice of the generator $\chi$.
(ii) We have the congruence
$\det A_{q}(k)\equiv(-1)^{\frac{m^{2}-m+2}{2}}\ ({\rm{mod}}\ p).$
The next theorem gives the explicit values of $\det A_{q}(1)$ and $\det A_{q}(2)$.
Theorem 1.2.
Let $q=p^{n}$ be a prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\chi$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then the following results hold.
(i) For the cyclotomic matrix $A_{q}(1)$, we have
$\det A_{q}(1)=(-1)^{\frac{(q-2)(q-3)}{2}}(q-1)^{q-1}.$
[…]
$\det A_{q}(2)=(-1)^{\alpha_{p}}\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}2^{\frac{p^{n-1}-1}{2}},$
where
$\alpha_{p}=\begin{cases}1&\mbox{if}\ n\equiv 1\ ({\rm{mod}}\ 2),\\ (p^{2}+7)/8&\mbox{if}\ n\equiv 0\ ({\rm{mod}}\ 2).\end{cases}$
Remark 1.1.
For an arbitrary positive integer $13\leq k<q-$ 1 with $k\mid q-1$, it seems that determining the explicit value of $\det A_{q}(k)$ is very difficult.
Theorem 1.3.
Let $q=p^{n}$ be a prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\chi$ be a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Then the following results hold.
(i) Let $k\mid p-1$ be a positive integer. Then $\det B_{p}(k)\neq 0$. In contrast, if $n\geq 2$, then we have $\det B_{q}(k)=0$.
(ii) For the case $k=1$ we have
$\det B_{p}(1)=\frac{(-1)^{\frac{p(p+1)}{2}}(p-1)^{p-1}}{p^{p-2}}.$
(iii) If $p$ is an odd prime, then
$\det B_{p}(2)=(-1)^{\frac{(p+3)(p-1)}{4}}p\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}.$
Remark 1.2.
For Theorem 1.3(i), the condition $k\mid p-1$ cannot be removed. In fact, with the help of computer we verify that
$\det B_{25}(3)\neq 0.$
Inspired by Theorem 1.3(i) and the above remark, we posed the following conjecture.
Conjecture 1.1.
Let $q=p^{n}$ be a prime power with $p$ prime and $n\geq 2$. Then for any positive integer $11\leq k\leq p-$ 1 with $k\mid q-1$ and $k\nmid p-1$ we have
$\det B_{q}(k)\neq 0.$
Remark 1.3.
We will give some detailed discussions on this conjecture in Section 5.
1.4. Outline of This Paper
We will prove Theorem 1.1 in Section 2. The proofs of Theorems 1.21.3 will be given in Sections 3–4 respectively. Finally, in Section 5 we will pose some problems for further research.
2. Proof of Theorem 1.1
Let $q=p^{n}$ be a prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $L=\mathbb{Q}(\zeta_{q-1},\zeta_{p})$ and let ${\mathcal{O}}_{L}$ be the ring of algebraic integers over $L$. Let $\mathfrak{p}$ be a prime ideal of ${\mathcal{O}}_{L}$ such that $p\in\mathfrak{p}$. By algebraic number theory, it is easy to verify that
${\mathcal{O}}_{L}/\mathfrak{p}\cong\mathbb{F}_{q}.$
Now let $\chi_{\mathfrak{p}}\in\widehat{\mathbb{F}_{q}^{\times}}$ be the Teichümuller character of $\mathfrak{p}$, i.e.,
$\chi_{\mathfrak{p}}(x\ {\rm mod}\ \mathfrak{p})\equiv x\ ({\rm{mod}}\ \mathfrak{p})$
for any $x\in{\mathcal{O}}_{L}$. One can verify that $\chi_{\mathfrak{p}}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Also, for any integer $20\leq r\leq q-$ 2, let
$r=\sum_{0\leq j\leq n-1}r_{j}p^{j}\ (0\leq r_{j}\leq p-1)$
be the decomposition of $r$ in base $p$. We define
$s(r):=\sum_{0\leq j\leq n-1}r_{j},$
and
$t(r):=\prod_{0\leq j\leq n-1}(r_{j}!).$
We begin with the well known Stickelberger congruence (see [5, Theorem 3.6.6]).
Lemma 2.1.
Let notations be as above. Then for any integer $20\leq r\leq q-$ 2 we have
$\frac{G_{q}(\chi_{\mathfrak{p}}^{-r})}{(\zeta_{p}-1)^{s(r)}}\equiv-\frac{1}{t(r)}\ ({\rm{mod}}\ \mathfrak{p}).$
Let $m$ be a positive integer. We next consider a permutation of $\mathbb{Z}/m\mathbb{Z}$. Let $a\in\mathbb{Z}$ with $\gcd(a,m)=1$. Then the map $x\ {\rm mod}\ m\mapsto ax\ {\rm mod}\ m$ induces a permutation $\tau_{a}(m)$ of $\mathbb{Z}/m\mathbb{Z}$. Lerch [9] determined the sign of this permutation.
Lemma 2.2.
Let notations be as above and let $\mathrm{sign}(\tau_{a}(m))$ denote the sign of $\tau_{m}(a)$. Then
$\mathrm{sign}(\tau_{m}(a))=\begin{cases}(\frac{a}{m})&\mbox{if}\ m\equiv 1\ ({\rm{mod}}\ 2),\\ 1&\mbox{if}\ m\equiv 2\ ({\rm{mod}}\ 4),\\ (-1)^{(a-1)/2}&\mbox{if}\ m\equiv 0\ ({\rm{mod}}\ 4),\end{cases}$
where $(\frac{\cdot}{m})$ is the Jacobi symbol if $m$ is odd. In particular, we have
(2.14) $\mathrm{sign}(\tau_{m}(-1))=(-1)^{\frac{(m-1)(m-2)}{2}}.$
For simplicity, $\sum_{a\in\mathbb{F}_{q}}$ and $\sum_{a\in\mathbb{F}_{q}^{\times}}$ will be abbreviated as $\sum_{a}$ and $\sum_{a\neq 0}$ respectively. Also, let $M_{m}({\mathcal{O}}_{L})$ denote the set of $m\times m$ matrices with all entries contained in ${\mathcal{O}}_{L}$. For any $M_{1},M_{2}\in M_{m}({\mathcal{O}}_{L})$, the symbol
$M_{1}\equiv M_{2}\ ({\rm{mod}}\ \mathfrak{p})$
means that every entry of $M_{1}-M_{2}$ is contained in the prime ideal $\mathfrak{p}$.
Now we are in a position to prove our first theorem.
Proof of Theorem 1.1. (i) Recall that $L=\mathbb{Q}(\zeta_{q-1},\zeta_{p})$. As $\gcd(q-1,p)=1$, we have $\mathbb{Q}(\zeta_{q-1})\cap\mathbb{Q}(\zeta_{p})=\mathbb{Q}$ and hence
${\rm Gal}(L/\mathbb{Q})\cong{\rm Gal}(L/\mathbb{Q}(\zeta_{q-1}))\oplus{\rm Gal}(L/\mathbb{Q}(\zeta_{p}))\cong\left(\mathbb{Z}/p\mathbb{Z}\right)^{\times}\oplus\left(\mathbb{Z}/(q-1)\mathbb{Z}\right)^{\times}.$
For any $\sigma_{l,s}\in{\rm Gal}(L/\mathbb{Q})$ with $l,s\in\mathbb{Z}$, $\gcd(l,p)=1$ and $\gcd(s,q-1)$, we have $\sigma_{l,s}(\zeta_{p})=\zeta_{p}^{l}$ and $\sigma_{l,s}(\zeta_{q-1})=\zeta_{q-1}^{s}$. Now for any $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$, one can verify that
$\sigma_{l,s}\left(G_{q}(\psi)\right)$ $=\sigma_{l,s}\left(\sum_{a}\psi(a)\zeta_{p}^{{\rm Tr}(a)}\right)$
$=\sum_{a}\psi^{s}(a)\zeta_{p}^{{\rm Tr}(la)}$
$=\psi^{-s}(l)\sum_{a}\psi^{s}(la)\zeta_{p}^{{\rm Tr}(la)}$
$=\psi^{-s}(l)\sum_{a}\psi^{s}(a)\zeta_{p}^{{\rm Tr}(a)}$
$=\psi^{-s}(l)G_{q}(\psi^{s}).$
Letting $m=(q-1)/k$, by the above and Lemma 2.2 we have
$\sigma_{l,s}\left(\det A_{q}(k)\right)$ $=\det\left[\chi^{-(ksi+ksj)}(l)G_{q}(\chi^{ksi+ksj})\right]_{0\leq i,j\leq m-1}$
$=\left(\prod_{0\leq r\leq m-1}\chi^{-ksr}(l)\right)^{2}\cdot\mathrm{sign}(\tau_{m}(s))^{2}\cdot\det A_{q}(k)$
$=\det A_{q}(k).$
Noting that $\gcd(s,m)=1$, the last equality follows from
$\left(\prod_{0\leq r\leq m-1}\chi^{-ksr}(l)\right)^{2}$ $=\prod_{0\leq r\leq m-1}\chi^{kr}(l)\cdot\prod_{0\leq r\leq m-1}\chi^{-kr}(l)=1.$
Also, for any integer $s$, the character $\chi^{s}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$ if and only if $\gcd(s,q-1)=1$. Hence for any generator $\chi^{\prime}$, there is an integer $s$ with $\gcd(s,q-1)=1$ such that
$\det\left[G_{q}(\chi^{\prime(ki+kj)})\right]_{0\leq i,j\leq m-1}=\sigma_{1,s}\left(\det A_{q}(k)\right)=\det A_{q}(k).$
In view of the above, we have proved (i) for $\det A_{q}(k)$. Using essentially the same method, one can also verify that (i) holds for $\det B_{q}(k)$.
(ii) By (2.14) of Lemma 2.2 we have
$\det A_{q}(k)=(-1)^{\frac{(m-1)(m-2)}{2}}\det\left[G_{q}(\chi^{ki-kj})\right]_{0\leq i,j\leq m-1}.$
By Lemma 2.1, for any $10\leq i\neq j\leq m-$ 1 we have
$G_{q}(\chi^{ki-kj})\equiv 0\ ({\rm{mod}}\ \mathfrak{p}).$
By this and noting that $G_{q}(\varepsilon)=-1$, we obtain
$\left[G_{q}(\chi^{ki-kj})\right]_{0\leq i,j\leq m-1}\equiv-I_{m}\ ({\rm{mod}}\ \mathfrak{p}),$
where $I_{m}$ is the $m\times m$ identity matrix. By the above, we have
$\det A_{q}(k)\equiv(-1)^{\frac{(m-1)(m-2)}{2}+m}\equiv(-1)^{\frac{m^{2}-m+2}{2}}\ ({\rm{mod}}\ p).$
This completes the proof of (ii).
In view of the above, we have completed the proof of Theorem 1.1.∎
3. Proof of Theorem 1.2
We begin with follow known result in linear algebra.
Lemma 3.3.
Let $m$ be a positive integer and let $M$ be an $m\times m$ complex matrix. Let $\lambda_{1},\cdots,\lambda_{m}\in\mathbb{C}$, and let ${\bm{v}}_{1},\cdots,{\bm{v}}_{m}\in\mathbb{C}^{m}$ be column vectors. Suppose that
$M{\bm{v}}_{i}=\lambda_{i}{\bm{v}}_{i}$
for each $1\leq i\leq m$ and that the vectors ${\bm{v}}_{1},\cdots,{\bm{v}}_{m}$ are linearly independent over $\mathbb{C}$. Then $\lambda_{1},\cdots,\lambda_{m}$ are exactly all the eigenvalues of $M$ (counting multiplicity).
Recall that $\chi$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$ and $k\mid q-1$ is a positive integer. Throughout the remaining part of this section, we let $m=(q-1)/k$. Let $U_{k}:=\{x\in\mathbb{F}_{q}:\ x^{k}=1\}$ be the set of all $m$th roots of unity in $\mathbb{F}_{q}$.
We also need the following known result.
Lemma 3.4.
Let notation be as above. Then for any $x\in\mathbb{F}_{q}$ we have
$\sum_{0\leq r\leq m-1}\chi^{kr}(x)=\begin{cases}m&\mbox{if}\ x\in U_{k},\\ 0&\mbox{otherwise.}\end{cases}$
Define
(3.15) $C_{q}(k):=\left[G_{q}(\chi^{ki-kj})\right]_{0\leq i,j\leq m-1}.$
The next lemma determines all the eigenvalues of $C_{q}(k)$.
Lemma 3.5.
Let notations be as above. For any element $b$ in the quotient group $\mathbb{F}_{q}^{\times}/U_{k}$, let
(3.16) $\lambda_{b}:=m\sum_{y\in U_{k}}\zeta_{p}^{{\rm Tr}(by)}.$
Then these $\lambda_{b}$ (where $b\in\mathbb{F}_{q}^{\times}/U_{k}$) are exactly all the eigenvalues of $C_{q}(k)$.
Proof.
For any $b\in\mathbb{F}_{q}^{\times}/U_{k}$, define the column vector
${\bm{v}}_{b}:=\left(\chi^{0}(b),\chi^{k}(b),\cdots,\chi^{k(m-1)}(b)\right)^{T}.$
Then by Lemma 3.4
$\sum_{0\leq r\leq m-1}G_{q}(\chi^{ki-kr})\chi^{kr}(b)$ $=\sum_{0\leq r\leq m-1}\sum_{a\neq 0}\chi^{ki}(a)\zeta_{p}^{{\rm Tr}(a)}\chi^{kr}\left(\frac{b}{a}\right)$
$=\sum_{a\neq 0}\chi^{ki}(a)\zeta_{p}^{{\rm Tr}(a)}\sum_{0\leq r\leq m-1}\chi^{kr}\left(\frac{b}{a}\right)$
$=m\sum_{y\in U_{k}}\chi^{ki}(by)\zeta_{p}^{{\rm Tr}(by)}$
$=\lambda_{b}\chi^{ki}(b).$
This implies $C_{q}(k){\bm{v}}_{b}=\lambda_{b}{\bm{v}}_{b}$ for each $b\in\mathbb{F}_{q}^{\times}/U_{k}$.
On the other hand, let
$V_{k}=\left[\chi^{ki}(b)\right]_{0\leq i\leq m-1,\ b\in\mathbb{F}_{q}^{\times}/U_{k}}.$
Then it is clear that
$(\det V_{k})^{2}=\pm\prod_{b,b^{\prime}\in\mathbb{F}_{q}^{\times}/U_{k},b\neq b^{\prime}}\left(\chi^{k}(b)-\chi^{k}(b^{\prime})\right)\neq 0.$
Hence these ${\bm{v}}_{b}$ are linearly independent over $\mathbb{C}$. Now by Lemma 3.3, these $\lambda_{b}$ (where $b\in\mathbb{F}_{q}^{\times}/U_{k}$) are exactly all the eigenvalues of $C_{q}(k)$. This completes the proof. ∎
We also need the following result concerning the trace map ${\rm Tr}:={\rm Tr}_{\mathbb{F}_{q}/\mathbb{F}_{p}}$ (see [5, Proposition 2.4.11]).
Lemma 3.6.
The trace map ${\rm Tr}$ is a surjective $\mathbb{F}_{p}$-homomorphism from $\mathbb{F}_{q}$ to $\mathbb{F}_{p}$. Also,
${\rm Ker}({\rm Tr})=\{x\in\mathbb{F}_{q}:\ {\rm Tr}(x)=0\}=\{x^{p}-x:\ x\in\mathbb{F}_{q}\}$
is an $\mathbb{F}_{p}$-subspace of dimension $n-1$.
Recall also that $L=\mathbb{Q}(\zeta_{p},\zeta_{q-1})$ and ${\mathcal{O}}_{L}$ is the ring of algebraic integers over $L$. Let ${\mathcal{O}}_{L}^{\times}$ be the ring of units in ${\mathcal{O}}_{L}$.
Now we are in a position to prove our second theorem.
Proof of Theorem 1.2. (i) By Lemma 2.2 and Lemma 3.5, one can verify that $\det A_{q}(1)$ is equal to
$\mathrm{sign}(\tau_{q-1}(-1))\cdot\det C_{q}(1)$ $=(-1)^{\frac{(q-2)(q-3)}{2}}(q-1)^{q-1}\prod_{b\in\mathbb{F}_{q}^{\times}}\zeta_{p}^{{\rm Tr}(b)}$
$=(-1)^{\frac{(q-2)(q-3)}{2}}(q-1)^{q-1}.$
The last equality follows from $\sum_{b\in\mathbb{F}_{q}^{\times}}{\rm Tr}(b)=0$.
(ii) Note that $2\nmid p$. We first consider $\det C_{q}(2)$ defined by (3.15). By Lemma 3.5 we have
(3.17) $\det C_{q}(2)=\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}\prod_{b\in\mathbb{F}_{q}^{\times}/U_{2}}\left(\zeta_{p}^{{\rm Tr}(b)}+\zeta_{p}^{{\rm Tr}(-b)}\right).$
Let
$\xi=\prod_{b\in\mathbb{F}_{q}^{\times}/U_{2}}\left(\zeta_{p}^{{\rm Tr}(b)}+\zeta_{p}^{{\rm Tr}(-b)}\right).$
Then by Lemma 2.2 and Theorem 1.1 clearly $\xi\in{\mathcal{O}}_{L}\cap\mathbb{Q}=\mathbb{Z}$. Hence by Lemma 3.6 we have
$\xi^{2}$ $=\prod_{b\in\mathbb{F}_{q}^{\times}/U_{2}}\left(\zeta_{p}^{{\rm Tr}(b)}+\zeta_{p}^{{\rm Tr}(-b)}\right)^{2}$
$=\prod_{b\in\mathbb{F}_{q}^{\times}}\left(\zeta_{p}^{{\rm Tr}(b)}+\zeta_{p}^{{\rm Tr}(-b)}\right)$
$=\prod_{b\in\mathbb{F}_{q}^{\times}}\zeta_{p}^{{\rm Tr}(-b)}\prod_{b\in\mathbb{F}_{q}^{\times}}\left(1+\zeta_{p}^{{\rm Tr}(2b)}\right)$
$=\prod_{b\in\mathbb{F}_{q}^{\times}}\left(1+\zeta_{p}^{{\rm Tr}(b)}\right)$
$=\prod_{b\in{\rm Ker}({\rm Tr})\setminus\{0\}}2\times\prod_{b\in\mathbb{F}_{q}^{\times}\setminus{\rm Ker}({\rm Tr})}\left(1+\zeta_{p}^{{\rm Tr}(b)}\right)$
$=2^{p^{n-1}-1}\prod_{b\in\mathbb{F}_{q}^{\times}\setminus{\rm Ker}({\rm Tr})}\left(1+\zeta_{p}^{{\rm Tr}(b)}\right).$
We claim that
$\mu:=\prod_{b\in\mathbb{F}_{q}^{\times}\setminus{\rm Ker}({\rm Tr})}\left(1+\zeta_{p}^{{\rm Tr}(b)}\right)=1.$
In fact, fix a positive integer $c$ with $2c\equiv 1\ ({\rm{mod}}\ p)$. Then for any $b\in\mathbb{F}_{q}^{\times}\setminus{\rm Ker}({\rm Tr})$ we have
$\left(1+\zeta_{p}^{{\rm Tr}(b)}\right)^{-1}=\frac{1-\zeta_{p}^{{\rm Tr}(b)}}{1-\zeta_{p}^{2{\rm Tr}(b)}}=\frac{1-\zeta_{p}^{2c{\rm Tr}(b)}}{1-\zeta_{p}^{2{\rm Tr}(b)}}\in{\mathcal{O}}_{L}.$
This implies $\mu\in{\mathcal{O}}_{L}^{\times}$. On the other hand, clearly $\mu=\xi^{2}/(2^{p^{n-1}-1})\in\mathbb{Q}$. Thus,
$\mu\in{\mathcal{O}}_{L}^{\times}\cap\mathbb{Q}=\{1,-1\}.$
As $\xi\in\mathbb{Z}$, we have $\mu=\xi^{2}/(2^{p^{n-1}-1})\geq 0$ and hence $\mu=1$.
In view of the above, we obtain
$\xi=\pm 2^{\frac{p^{n-1}-1}{2}}$
and hence by (3.17) we have
$\det C_{q}(2)=\pm\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}\cdot 2^{\frac{p^{n-1}-1}{2}}.$
By Lemma 2.2 we also have
$\det A_{q}(2)=\pm\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}\cdot 2^{\frac{p^{n-1}-1}{2}}.$
Recall that $q=p^{n}$. Since $p>2$, it is easy to verify that
$\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}\cdot 2^{\frac{p^{n-1}-1}{2}}\equiv(-1)^{\frac{q-1}{2}}\left(\frac{2}{p}\right)\equiv(-1)^{\frac{q-1}{2}+\frac{p^{2}-1}{8}}\ ({\rm{mod}}\ p).$
Combining the above results with Theorem 1.1(ii), we finally obtain
$\det A_{q}(2)=(-1)^{\alpha_{p}}\left(\frac{q-1}{2}\right)^{\frac{q-1}{2}}2^{\frac{p^{n-1}-1}{2}},$
where
$\alpha_{p}=\begin{cases}1&\mbox{if}\ n\equiv 1\ ({\rm{mod}}\ 2),\\ (p^{2}+7)/8&\mbox{if}\ n\equiv 0\ ({\rm{mod}}\ 2).\end{cases}$
This completes the proof of Theorem 1.2. ∎
4. Proof of Theorem 1.3
Recall that $q=p^{n}$ and $k\mid q-1$ is a positive integer. In this section, we also let $m=(q-1)/k$. We begin with the following known result (see [2, Theorem 1.1.4]).
Lemma 4.7.
Let $\psi\in\widehat{\mathbb{F}_{q}^{\times}}$ be a nontrivial character. Then
$G_{q}(\psi)G_{q}(\psi^{-1})=\psi(-1)q.$
Using Lemma 4.7, it is easy to verify the following result.
Lemma 4.8.
Let notations be as above. Then
$\left[\frac{1}{G_{q}(\chi^{-ki+kj})}\right]_{0\leq i,j\leq m-1}=\frac{1-q}{q}I_{m}+\frac{1}{q}D_{q}(k),$
where
(4.18) $D_{q}(k):=\left[(-1)^{ki-kj}G_{q}(\chi^{ki-kj})\right]_{0\leq i,j\leq m-1}.$
Moreover, we have
$P_{k}^{-1}C_{q}(k)P_{k}=D_{q}(k),$
where $P_{k}$ is the diagonal matrix
${\rm diag}\left((-1)^{k},(-1)^{2k},\cdots,(-1)^{(m-1)k}\right),$
and $C_{q}(k)$ is defined by (3.15). Hence $D_{q}(k)$ and $C_{q}(k)$ have the same eigenvalues.
We also need the following result due to Newman [10, Theorem 1].
Lemma 4.9.
Let $N\geq 2$ be a positive integer and let $\zeta_{N}=e^{2\pi{\bf i}/N}$. Let
$\Omega_{N}=\left\{l\in\mathbb{Z}^{+}:\exists\ c_{1},\cdots,c_{l}\in\mathbb{Z}^{+}\ \text{and}\ 0\leq e_{1}<\cdots<e_{l}\leq n-1\ \text{such that}\ \sum_{j=1}^{l}c_{j}\zeta_{N}^{e_{j}}=0\right\}.$
Then $\min(\Omega_{N})=p(N)$, where $p(N)$ is the least prime divisor of $N$.
Proof of Theorem 1.3. (i) Suppose first $n=1$. The case $k=p-1$ is trivial. Now we consider the case $11\leq k<p-$ 1. Then by Lemma 4.8, every eigenvalue of $[G_{p}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$ can be written as
$\theta_{b}:=\frac{1-p}{p}+\frac{m}{p}\sum_{y\in U_{k}}\zeta_{p}^{by}=\frac{m}{p}\sum_{y\in U_{k}}\left(\zeta_{p}^{by}-1\right)$
where $b\in\mathbb{F}_{p}^{\times}/U_{k}$. As $k<p-1$, by Lemma 4.9 we have
$\theta_{b}=\frac{m}{p}\left(-k\zeta_{p}^{0}+\sum_{y\in U_{k}}\zeta_{p}^{by}\right)\neq 0.$
This implies that $[G_{p}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$ is nonsingular. Now by Lemma 2.2 we have
$\det B_{p}(k)=\pm\det\left[\frac{1}{G_{p}(\chi^{-ki+kj})}\right]_{0\leq i,j\leq m-1}\neq 0.$
Suppose now $n\geq 2$. As $k\mid p-1$, we have $U_{k}\subseteq\mathbb{F}_{p}^{\times}$. By Lemma 3.6 there is an element $b\in\mathbb{F}_{q}^{\times}$ with ${\rm Tr}(b)=0$. Similar to the above, we see that
$\theta_{b}$ $=\frac{1-q}{q}+\frac{m}{q}\sum_{y\in U_{k}\subseteq\mathbb{F}_{p}^{\times}}\zeta_{p}^{{\rm Tr}(by)}$
$=\frac{1-q}{q}+\frac{m}{q}\sum_{y\in U_{k}\subseteq\mathbb{F}_{p}^{\times}}\zeta_{p}^{y{\rm Tr}(b)}$
$=0$
is an eigenvalue of $[G_{q}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$. By Lemma 2.2 again, $B_{q}(k)$ is singular.
(ii) Observe first that
(4.19) $\prod_{b\in\mathbb{F}_{p}^{\times}}\left(1-\zeta_{p}^{b}\right)=\lim_{x\rightarrow 1}\frac{x^{p}-1}{x-1}=p.$
Using this and by Lemma 4.8, Lemma 3.5 and (3.16), we have
$\det[G_{p}(\chi^{-i+j})^{-1}]_{0\leq i,j\leq p-2}$ $=\prod_{b\in\mathbb{F}_{p}^{\times}}\left(\frac{1-p}{p}+\frac{p-1}{p}\zeta_{p}^{b}\right)$
$=\left(\frac{1-p}{p}\right)^{p-1}\prod_{b\in\mathbb{F}_{p}^{\times}}(1-\zeta_{p}^{b})$
$=\left(\frac{1-p}{p}\right)^{p-1}p.$
By Lemma 2.2 we have
$\det B_{p}(1)=\frac{(-1)^{\frac{p(p+1)}{2}}(p-1)^{p-1}}{p^{p-2}}.$
(iii) By Lemma 4.8, Lemma 3.5, (3.16) and (4.19) again, and noting that $2\nmid p$, one can verify that
$\det[G_{p}(\chi^{-2i+2j})^{-1}]_{0\leq i,j\leq(p-3)/2}$ $=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}/U_{2}}\left(\zeta_{p}^{b}+\zeta_{p}^{-b}-2\right)$
$=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}/U_{2}}\left(\zeta_{p}^{2b}+\zeta_{p}^{-2b}-2\right)$
$=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}/U_{2}}\left(\zeta_{p}^{b}-\zeta_{p}^{-b}\right)^{2}$
$=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\cdot(-1)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}}\left(\zeta_{p}^{b}-\zeta_{p}^{-b}\right)$
$=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\cdot(-1)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}}\zeta_{p}^{b}\left(1-\zeta_{p}^{-2b}\right)$
$=\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}\cdot(-1)^{\frac{p-1}{2}}\prod_{b\in\mathbb{F}_{p}^{\times}}\left(1-\zeta_{p}^{b}\right)$
$=(-1)^{\frac{p-1}{2}}p\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}.$
Now by Lemma 2.2 we obtain
$\det B_{p}(2)=(-1)^{\frac{(p+3)(p-1)}{4}}p\left(\frac{p-1}{2p}\right)^{\frac{p-1}{2}}.$
In view of the above, we have completed the proof of Theorem 1.3.∎
5. Concluding Remarks
In this section we focus on Conjecture 1.1. By Theorem 1.3 and the previous proofs, it is easy to see that $\det B_{q}(k)=0$ if and only if there is an element $b\in\mathbb{F}_{q}^{\times}$ such that
(5.20) $\sum_{y\in U_{k}}\left(\zeta_{p}^{{\rm Tr}(by)}-1\right)=\left(-k\zeta_{p}^{0}+\sum_{y\in U_{k}}\zeta_{p}^{{\rm Tr}(by)}\right)=0.$
As $11\leq k<p-$ 1, by Lemma 4.9 we see that (5.20) holds if and only if ${\rm Tr}(by)=0$ for any $y\in U_{k}$.
Set $U_{k}=\{y_{1},\cdots,y_{k}\}$. By the above discussion, Conjecture 1.1 is equivalent to saying that the system of equations
$\begin{cases}y_{1}x+y_{1}^{p}x^{p}+\cdots+y_{1}^{p^{n-1}}x^{p^{n-1}}=0\\ y_{2}x+y_{2}^{p}x^{p}+\cdots+y_{2}^{p^{n-1}}x^{p^{n-1}}=0\\ \cdots\\ y_{k}x+y_{k}^{p}x^{p}+\cdots+y_{k}^{p^{n-1}}x^{p^{n-1}}=0\end{cases}$
has no nonzero solutions over $\mathbb{F}_{q}$. However, we cannot solve this problem currently.
Acknowledgments We would like to thank Prof. Hao Pan for his steadfast encouragement. This work was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12201291).